# DepsRAG: Towards Managing Software Dependencies using Large Language Models

### v2

@@ 1. Introduction
Today, developers are expected to manage their software dependencies (Kula et al., 2018). However, in a recent survey, developers reported that they are not confident in their current dependency management practices (Tidelift, [n. d.]). Although developers often have full control over which dependency packages and versions to use, they have no control over the dependencies of the packages used. Such transitive dependencies, which often amount to more than direct dependencies, may have various vulnerabilities that put software projects at risk. To mitigate these risks, we developed DepsRAG to support practitioners and software developers understand the dependencies in their software supply chain. DepsRAG facilitates constructing and analyzing software dependencies for four popular software ecosystems. It also answers key questions about the dependency graph, including the number of dependencies, the depth of the dependency graph, and the number of paths/chains. The use of LLM enables DepsRAG to autonomously generate queries for KGs, unlike prior methods demanding hardcoded or user-generated queries (Dann et al., 2023; Litzenberger et al., 2023; Yamaguchi et al., 2014), thereby enhancing scalability. This retrieved information enhances the user’s question, referred to as Retrieval Augmented Generation (RAG), enabling accurate LLM responses. DepsRAG works in two phases. In the first phase, the user provides the name and ecosystem of the project. Then it transitively constructs the dependencies as a KG. In the second phase, the user can chat with DepsRAG to answer questions about the dependencies. DepsRAG is equipped with several tools to facilitate its functionality to access the KG, automatically generate queries, search the Web, and visualize the dependency graph.
@@ 2.1. Dependency Graphs
Dependency graphs are a fundamental concept in software engineering and are used to represent dependencies among various entities of a software system. These entities can be classes, functions, modules, packages, or even larger components. Each node in the graph represents an entity, and each directed edge between two nodes indicates that one entity depends on another. This means that changes in one entity can affect the other.
Litzenberger et al. (Litzenberger et al., 2023) proposed a framework the defines three levels of granularity for dependency representation: package-to-package, artifact-to-package, and artifact-to-artifact. Düsing and Hermann (Düsing and Hermann, 2022) constructed a dependency graph at the artifact-to-package level and leveraged these graphs to investigate the propagation of vulnerabilities in different software repositories. Benelallam et al. (Benelallam et al., 2019) constructed the Maven Dependency Graph to answer high-level research questions about artifact releases, evolution, and usage trends over time. Dependency graphs have also been used to build tools for program understanding (Falke et al., 2005; Musco et al., 2017).
@@ 2.3. Retrieval-Augmented Generation (RAG)
RAG (Lewis et al., 2020; Guu et al., 2020) augments the input questions for LLMs with information retrieved from documents, knowledge graphs, or vector databases, and is considered a useful method for addressing LLM hallucinations. As such, RAG provides extra knowledge sources for LLMs and greatly improves their performance in a variety of tasks, especially in the knowledge-intensive ones. RAG applications leverage tools and function calls to perform their job. This feature is introduced by LLM multi-agent frameworks (i.e., LangChain (Chase, 2022) and Langroid (lan, 2024)) and recent versions of OpenAI models. [^2] https://platform.openai.com/docs/guides/text-generation/function-calling This feature enables users to provide function descriptions to the LLM. The LLM responds with a structured response, such as a JSON file containing the required arguments to invoke any available functions. These functions serve as action executors and provide the option to supply APIs that the LLM can query to obtain essential information for responding to users. For example, when a user asks about the current weather in a specific city and provides the LLM with a weather API call description, the LLM can augment its response with information retrieved from the API. This facilitates a seamless and efficient interaction between the user and the LLM and provides rich responses that avoid hallucinations.
@@ 3. Approach
  • Web Search Retriever. This component acts as a failover mechanism and answers when no answer is obtained from the KG. It also provides answers beyond the scope of the KG. For example, the user could ask DepsRAG if one of the packages in the dependency KG is vulnerable. This component leverages WebSearchTool to perform a Web search and retrieve top-3 results, and then the LLM can extract the most relevant results.
Figure 3 outlines the prompt used by DepsRAG. This prompt consists of three parts. R defines the role of DepsRAG. $\textrm{p}_{1}$ and $\textrm{p}_{2}$ describe the instructions for the two phases of DepsRAG:
@@
    Dependency KG Construction. The user provides the name of the package, the version, and the ecosystem, and then DepsRAG constructs the dependency graph following the graph schema illustrated in Figure 2. According to this schema, the nodes in the KG belong to the entity Package, which has only two properties, the package name and version. For entity relationships, there is only one relation depends_on. DepsRAG leverages ConstructKGTool for constructing the KG. Suppose the user did not provide the version and ecosystem of the package. In this case, DepsRAG triggers the Web Search Retriever to find the version and ecosystem based on the provided package name. Figure 3 illustrates the workflow of this phase to construct the dependency graph of the OpenAI package in the PyPI ecosystem.
@@
Figure 3: DepsRAG Instructions R Define LLM Role: You are an expert in software dependency graphs and analyzing them using Neo4j.
$\textrm{p}_{1}$ Phase-1 Task Instructions: FIRST, I’ll give you the name of the package that I want to analyze.
THEN, you can use the ”WebSearchTool” to find out information about a package, such as version number and package type (PyPI, NPM, Cargo, or GO). If you cannot get this information, ask me and I can tell you.
DON’T forget to include the package name in your questions. After receiving this information, ensure to get the package version type. THEN ask the user if they want to construct the dependency graph, and if so, use ”ConstructKGTool” to construct the dependency graph. Otherwise, say ”Could not retrieve package type or version” and NO_ANSWER.
@@
Figure 2. Dependency KG Schema consisting of a single entity with two attributes, name and version, and one relation.
@@ 5. Related Work
Litzenberger et al. (Litzenberger et al., 2023) proposed a unified data model to implement and construct dependency graphs for arbitrary repositories, thus facilitating the comparison of dependency graphs between different repositories. In their implementation, the data model is based on Neo4j, which makes is compliant with DepsRAG’s KG schema. Maninger et al. (Maninger et al., 2023) proposed a visionary approach involving KGs to create a trustworthy AI software development assistant. Specifically, KGs can enable LLMs to correctly and appropriately explain the generated code. To our knowledge, we are the first to investigate the ability of LLMs to generate KGs for software dependencies, and utilize these graphs to aid in responding to queries concerning the dependency structure.

### v3

@@ 1. Introduction
Today, developers are expected to manage their software dependencies (Kula et al., 2018). However, in a recent survey, developers reported that they are not confident in their current dependency management practices (Tidelift, [n. d.]). Although developers often have full control over which dependency packages and versions to use, they have no control over the dependencies of the packages used. Such transitive dependencies, which often amount to more than direct dependencies, may have various vulnerabilities that put software projects at risk. To mitigate these risks, we developed DepsRAG to support practitioners and software developers in understanding the dependencies in their software supply chain. DepsRAG enables the construction and analysis of software dependencies for four popular ecosystems, answering key questions about the dependency graph, such as the number of dependencies, graph depth, and number of paths. Using LLM, DepsRAG autonomously generates queries for KGs, unlike previous methods requiring hardcoded or user-generated queries (Dann et al., 2023; Litzenberger et al., 2023; Yamaguchi et al., 2014), enhancing scalability. This retrieved information is used to augment the users’ questions, known as Retrieval Augmented Generation (RAG), which improves the accuracy of LLM responses. DepsRAG works in two phases. In the first phase, the user provides the name and ecosystem of the project. Then it transitively constructs the dependencies as a KG. In the second phase, the user can chat with DepsRAG to answer questions about the dependencies. DepsRAG is equipped with several tools to facilitate its functionality to access the KG, automatically generate queries, search the Web, and visualize the dependency graph.
Value. Managing software dependencies has become crucial for ensuring compliance with regulations like the EU Cyber Resilience Act (CRA) (CRA, [n. d.]) and the US Executive Order (EO) 14028 (nis, [n. d.]). It is also a key part of vulnerability management programs. Thus, there is a significant need to equip practitioners with effective tools that can operate autonomously at scale, without requiring deep expertise.
Impact. The application of LLMs to manage software dependencies offers usability, scalability, and autonomy, while facilitating seamless integration with other tools and retrievers, as opposed to the limited scope of ad-hoc manual dependency analysis. Representing dependencies as KGs provides a robust foundation to leverage existing sources of software dependencies. As a result, these features address the need for a systematic approach to consuming documents that record dependencies (i.e., software bill of materials) (Xia et al., 2023), and encourage researchers to investigate this approach to resolve various software supply chain issues.
@@ 2.1. Dependency Graphs
Dependency graphs are essential in software engineering for representing dependencies among software entities like classes, functions, modules, packages, or larger components. Each node represents an entity and directed edges indicate dependencies, meaning changes in one entity can affect another.
Litzenberger et al. (Litzenberger et al., 2023) proposed a framework with three levels of dependency granularity: package-to-package, artifact-to-package, and artifact-to-artifact. Düsing and Hermann (Düsing and Hermann, 2022) used artifact-to-package graphs to study vulnerability propagation in software repositories. Benelallam et al. (Benelallam et al., 2019) created the Maven Dependency Graph to explore artifact releases, evolution, and usage trends. Dependency graphs are also utilized in tools for program understanding (Falke et al., 2005; Musco et al., 2017).
@@ 2.3. Retrieval-Augmented Generation (RAG)
RAG (Lewis et al., 2020; Guu et al., 2020) augments LLM input questions with information from documents, knowledge graphs, or vector databases, effectively addressing LLM hallucinations. This method enhances LLM performance, particularly in knowledge-intensive tasks. RAG applications utilize tools and function calls, a feature introduced by LLM multi-agent frameworks like LangChain (Chase, 2022) and Langroid (lan, 2024), as well as recent OpenAI models. [^2] https://platform.openai.com/docs/guides/text-generation/function-calling This feature allows users to provide function descriptions to the LLM, which responds with structured outputs like JSON files containing necessary arguments to invoke functions. These functions can query APIs for essential information, enabling efficient, rich responses that avoid hallucinations. For example, the LLM can use a weather API to provide current weather information when asked.
@@ 3. Approach
  • Web Search Retriever. This component acts as a failover mechanism and answers when no answer is obtained from the KG. It also provides answers beyond the scope of the KG. For example, the user could ask DepsRAG if one of the packages in the dependency KG is vulnerable. This component leverages WebSearchTool to perform a Web search and retrieve top-3 outcomes, and then the LLM can extract the most relevant results.
Prompt 3 outlines the prompt used by DepsRAG. This prompt consists of three parts. R defines the role of DepsRAG. $\textrm{p}_{1}$ and $\textrm{p}_{2}$ describe the instructions for the two phases of DepsRAG:
@@
    Dependency KG Construction. The user provides the name of the package, the version, and the ecosystem, and then DepsRAG constructs the dependency graph following the graph schema illustrated in Figure 2. According to this schema, the nodes in the KG belong to the entity Package, which has only two properties, the package name and version. For entity relationships, there is only one relation “depends_on”. DepsRAG leverages ConstructKGTool for constructing the KG. Suppose the user did not provide the version and ecosystem of the package. In this case, DepsRAG triggers the Web Search Retriever to find the version and ecosystem based on the provided package name. Figure 3 illustrates the workflow of this phase to construct the dependency graph of the OpenAI package in the PyPI ecosystem.
@@
Prompt 1: DepsRAG Instructions R Define LLM Role: You are an expert in software dependency graphs and analyzing them using Neo4j.
$\textrm{p}_{1}$ Phase-1 Task Instructions: FIRST, I’ll give you the name of the package that I want to analyze. THEN, you can use the ”WebSearchTool” to find out information about a package, such as version number and package type (PyPI, NPM, Cargo, or GO). If you cannot get this information, ask me and I can tell you. DON’T forget to include the package name in your questions. After receiving this information, ensure to get the package version type. THEN ask the user if they want to construct the dependency graph, and if so, use ”ConstructKGTool” to construct the dependency graph. Otherwise, say ”Could not retrieve package type or version” and NO_ANSWER.
@@
Figure 2. Dependency KG Schema consisting of a single entity with two attributes, “name” and “version”, and one relation.
@@ 5. Related Work
Litzenberger et al. (Litzenberger et al., 2023) proposed a unified data model to implement and construct dependency graphs for arbitrary repositories, thus facilitating the comparison of dependency graphs between different repositories. In their implementation, the data model is based on Neo4j, which makes is compliant with DepsRAG’s KG schema. Maninger et al. (Maninger et al., 2024) proposed a visionary approach involving KGs to create a trustworthy AI software development assistant. Specifically, KGs can enable LLMs to correctly and appropriately explain the generated code. To our knowledge, we are the first to investigate the ability of LLMs to generate KGs for software dependencies, and utilize these graphs to aid in responding to queries concerning the dependency structure.